### v1

# A note on separability in outer automorphism groups

Abstract
We give a criterion for separability of subgroups of certain outer automorphism groups. This answers questions of Hagen and Sisto, by strengthening and generalizing a result of theirs on mapping class groups.
1 Introduction
A subgroup $H\leqslant G$ is called separable if for all $g\in G\setminus H$ there exists a finite quotient $q\colon G\to Q$ such that $q(g)\notin q(H)$. A question of Reid [Rei06, Question 3.5] asks whether all convex-cocompact subgroups of mapping class groups (as defined in [FM02], see also [KL08] for a characterization) are separable. Some examples were given in [LM07]. Hagen and Sisto show in [HS23] that certain examples of free convex-cocompact subgroups, constructed by Mj [Mj], are separable. In order to do so, they prove the following criterion.
Theorem ([HS23, Theorem 1.1]).
Let $g\geqslant 2$, and let $H\leqslant\operatorname{MCG}(\Sigma_{g})$. Suppose that $H$ is torsion-free, malnormal, and convex-cocompact. If the preimage of $H$ in $\operatorname{MCG}(\Sigma_{g}\setminus\{p\})$, for some point $p\in\Sigma_{g}$, is conjugacy separable, then $H$ is separable.
Remark.
The theorem is stated for $g\geqslant 1$, however in that case $\operatorname{MCG}(\Sigma_{g})$ is virtually free so every finitely generated subgroup is separable. We focus on the case $g\geqslant 2$ for coherence with the results of this note.
A group $G$ is said to be conjugacy separable if for all non-conjugate elements $g,h\in G$ there exists a finite quotient $q\colon G\to Q$ such that $q(g)$ and $q(h)$ are non-conjugate. Our first result removes the other hypotheses on $H$, answering [HS23, Question 1.4].
Theorem A.
Let $g\geqslant 2$, and let $H\leqslant\operatorname{MCG}(\Sigma_{g})$. If the preimage of $H$ in $\operatorname{MCG}(\Sigma_{g}\setminus\{p\})$, for some point $p\in\Sigma_{g}$, is conjugacy separable, then $H$ is separable.
We can replace the hypothesis of conjugacy separability by a more geometric one, which goes in the direction of Reid’s question [Rei06, Question 3.5]. This is the same way that Hagen and Sisto verify that their criterion holds for the groups constructed by Mj.
Corollary B.
Let $g\geqslant 2$, and let $H\leqslant\operatorname{MCG}(\Sigma_{g})$ be a convex-cocompact subgroup. If the preimage of $H$ in $\operatorname{MCG}(\Sigma_{g}\setminus\{p\})$, for some point $p\in\Sigma_{g}$, acts properly and cocompactly on a $\operatorname{CAT(0)}$ cube complex, then $H$ is separable.
Theorem A will follow from a more general result.
Theorem C.
Let $G$ be a finitely generated group with trivial center, and let $H\leqslant\operatorname{Out}(G)$. Suppose that $\operatorname{Aut}(G)$ is acylindrically hyperbolic and has no non-trivial finite normal subgroups. If the preimage of $H$ in $\operatorname{Aut}(G)$ is conjugacy separable, then $H$ is separable.
Thanks to recent works proving acylindrical hyperbolicity of automorphism groups [Gen19, GH21, GM19, Gen, EH], this can be applied in several contexts. We isolate two instances.
Corollary D.
Let $G$ be a torsion-free non-elementary hyperbolic group, and let $H\leqslant\operatorname{Out}(G)$. If the preimage of $H$ in $\operatorname{Aut}(G)$ is conjugacy separable, then $H$ is separable.
Corollary E.
Let $\Gamma$ be a finite simplicial graph with more than one vertex that does not decompose as a join of two non-empty subgraphs. Let $A_{\Gamma}$ be the corresponding right-angled Artin group, and let $H\leqslant\operatorname{Out}(A_{\Gamma})$. If the preimage of $H$ in $\operatorname{Aut}(A_{\Gamma})$ is conjugacy separable, then $H$ is separable.
Both of these corollaries apply to $\operatorname{Out}(F_{n})$. In this case, there are notions of convex-cocompact subgroups [DT18, HH18, DT17]; the corresponding preimage in $\operatorname{Aut}(F_{n})$ is hyperbolic [DT18], but we do not know of instances in which conjugacy separability of such groups is known (besides the case of cyclic subgroups). We remark that free-by-cyclic groups $F_{n}\rtimes\mathbb{Z}$ often have non-separable subgroups [Kud24].
The proof of Theorem C involves an element of $G$ that recognizes non-inner automorphisms. The existence of such an element relies on the acylindrical hyperbolicity of $\operatorname{Aut}(G)$, and is what allows to strengthen and generalize the criterion of Hagen and Sisto, with a shorter proof. It partially answers [HS23, Question 1.5].
Acknowledgements.
The author thanks Alessandro Sisto, Ric Wade and Henry Wilton for useful comments.
2 Proofs
We will use $\pi$ to denote the homomorphism $\pi\colon\operatorname{Aut}(G)\to\operatorname{Out}(G)$, and for a subgroup $H\leqslant\operatorname{Out}(G)$ we denote $\widetilde{H}\coloneqq\pi^{-1}H\leqslant\operatorname{Aut}(G)$. The starting point is the following criterion for separability, which in turn is based on Grossman’s criterion for residual finiteness [Gro75].
Proposition 2.1 ([HS23, Proposition 2.5]).
Let $G$ be a finitely generated group with trivial center. Let $H\leqslant\operatorname{Out}(G)$ and $\alpha\in\operatorname{Aut}(G)$. Suppose that:
  1. 1.
    There exists $x\in G$ such that $\alpha(x)\neq h(x)$ for all $h\in\widetilde{H}$;
  2. 2.
    $\widetilde{H}$ is conjugacy separable.
Then there exists a finite quotient $q\colon\operatorname{Aut}(G)\to Q$ such that $q(\alpha)\notin q(H)$.
Let us formulate the special case that we will use:
Corollary 2.2.
Let $G$ be a finitely generated group with trivial center and let $H\leqslant\operatorname{Out}(G)$. Suppose that:
  1. 1.
    There exists $\gamma\in\operatorname{Inn}(G)$ such that the centralizer of $\gamma$ in $\operatorname{Aut}(G)$ is $\langle\gamma\rangle$;
  2. 2.
    $\widetilde{H}$ is conjugacy separable.
Then $H$ is separable.
Proof.
We show that Proposition 2.1 holds for $H$ and an arbitrary $\alpha\in\operatorname{Out}(G)\setminus\widetilde{H}$. Choose $x\in G$ such that the corresponding inner automorphism $\gamma_{x}$ is as in the first assumption of the corollary. Suppose that $\alpha(x)=h(x)$ for some $h\in\widetilde{H}$. Then $h^{-1}\alpha(x)=x$, so $h^{-1}\alpha\in\langle\gamma_{x}\rangle\leqslant\operatorname{Inn}(G)\leqslant\widetilde{H}$, and which contradicts $\alpha\notin\widetilde{H}$. ∎
Acylindrical hyperbolicity of $\operatorname{Aut}(G)$ ensures that the first item holds.
Lemma 2.3.
Let $G$ be a group such that $\operatorname{Aut}(G)$ is acylindrically hyperbolic and has no non-trivial finite normal subgroups. Then there exists $\gamma\in\operatorname{Inn}(G)$ such that the centralizer of $\gamma$ in $\operatorname{Aut}(G)$ is $\langle\gamma\rangle$.
Proof.
We fix a hyperbolic Cayley graph $Y$ of $\operatorname{Aut}(G)$ such that the action is acylindrical and non-elementary [Osi16, Theorem 1.2]. In an acylindrically hyperbolic group every infinite normal subgroups is non-elementary [Osi16, Lemma 7.2]. In particular, $\operatorname{Inn}(G)\leqslant\operatorname{Aut}(G)$ is non-elementary (note that the assumptions imply that $G\cong\operatorname{Inn}(G)$ is infinite). Therefore there exists an inner automorphism $\gamma\in\operatorname{Inn}(G)$ that is loxodromic on $Y$, such that the elementary closure of $\gamma$ in $\operatorname{Aut}(G)$ is reduced to $\langle\gamma\rangle$ [Hul16, Lemma 5.6]. In particular, the centralizer of $\gamma$ in $\operatorname{Aut}(G)$ is reduced to $\langle\gamma\rangle$ [DGO17, Corollary 6.6]. ∎
Remark 2.4.
The same arguments give an element $x\in G$ such that, if $\alpha\in\operatorname{Aut}(G)$ is such that $\alpha(x)$ is conjugate to $x$, then $\alpha$ is inner. This gives a strong answer to [HS23, Question 1.5] under the same assumptions as Theorem C. Hagen and Sisto ask for a weaker property in all acylindrically hyperbolic groups - it is an open question whether the automorphism group of a finitely generated acylindrically hyperbolic group is always acylindrically hyperbolic [Gen19, Question 1.1]. They isolate the case of torsion-free non-elementary hyperbolic groups, which is covered by this argument, see the proof of Corollary D.
Proof of Theorem C.
Combine Lemma 2.3 and Corollary 2.2. ∎
For the next applications, we will use the following criterion to check that an automorphism group has no non-trivial finite normal subgroups. We say that $G$ has the unique root property if $x^{n}=y^{n}$ for some $x,y\in G,n\geqslant 1$ implies $x=y$.
Lemma 2.5.
Let $G$ be a group with trivial center and with the unique root property. Then $\operatorname{Aut}(G)$ has no non-trivial finite normal subgroups.
Proof.
Suppose that $N\leqslant\operatorname{Aut}(G)$ is a finite normal subgroup. The action of $G\cong\operatorname{Inn}(G)$ on $N$ by conjugacy has a finite index kernel $K$, that is $N$ fixes each element of $K$. Now let $\alpha\in N$ and $x\in G$. Let $n\geqslant 1$ be such that $x^{n}\in K$. Then $x^{n}=\alpha(x^{n})=\alpha(x)^{n}$, so by the unique root property $\alpha(x)=x$. This shows that $\alpha$ is trivial and we conclude. ∎
Proof of Corollary D.
If $G$ is torsion-free hyperbolic, then $\operatorname{Aut}(G)$ is acylindrically hyperbolic [GH21, Theorem 1.3]. Moreover, $G$ has trivial center and the unique root property [BB10, Lemma 2.2] and so $\operatorname{Aut}(G)$ has no non-trivial finite normal subgroups by Lemma 2.5. Therefore Theorem C applies. ∎
Proof of Corollary E.
If $G=A_{\Gamma}$ is as in the statement, then $\operatorname{Aut}(G)$ is acylindrically hyperbolic [Gen, Theorem 1.5]. Moreover, $G$ has trivial center and the unique root property [DK93, 3-2) and 3-3)] and so $\operatorname{Aut}(G)$ has no non-trivial finite normal subgroups by Lemma 2.5. Therefore Theorem C applies. ∎
We end by applying Corollary D to the special case of surface groups.
Proof of Theorem A.
Recall the Dehn–Nielsen–Baer Theorem: $\operatorname{MCG}(\Sigma_{g})$ is an index-$2$ subgroup of $\operatorname{Out}(\pi_{1}(\Sigma_{g}))$, and $\operatorname{MCG}(\Sigma_{g}\setminus\{p\})$ is the corresponding index-$2$ subgroup of $\operatorname{Aut}(\pi_{1}(\Sigma_{g}))$. Given a subgroup $H\leqslant\operatorname{MCG}(\Sigma_{g})$, its preimage in $\operatorname{MCG}(\Sigma_{g}\setminus\{p\})$ is the same as its preimage in $\operatorname{Aut}(\pi_{1}(\Sigma_{g}))$. Since $\pi_{1}(\Sigma_{g})$ is torsion-free non-elementary hyperbolic, we can apply Corollary D to get separability of $H$ in $\operatorname{Out}(\pi_{1}(\Sigma_{g}))$, which then implies separability in $\operatorname{MCG}(\Sigma_{g})$. ∎
Proof of Corollary B.
By Theorem A, it suffices to show that, under the hypotheses, the preimage $\widetilde{H}$ is conjugacy separable. By convex-cocompactness of $H$, $\widetilde{H}$ is hyperbolic [FM02, Ham]. By assumption $\widetilde{H}$ acts properly and cocompactly on a $\operatorname{CAT(0)}$ cube complex, and so it is virtually compact special [Ago13, Wis21]. It follows that $\widetilde{H}$ is conjugacy separable [MZ16]. ∎